# A Multi-scale Information Integration Framework for Infrared and Visible Image Fusion

### v1

@@ IV Experiments
In this section, we verify the effectiveness of our method through qualitative and quantitative evaluations by comparing it with several state-of-the-art methods on the TNO and RoadScene datasets. These comparative methods include the CNN-based method, GAN-based methods and AE-based methods, which are DenseFuse [12], RFN-Nest [23], GANMcC [15], STDFusionNet [40], DRF [38], UMF-CMGR[31] and YDTR[30]. Qualitiative evaluation is based on the human visual system, hoping to highlight the thermal target throughout the scene while preserving texture detail. Quantitative evaluation access fusion image quality the way of several statistics and visual metrics. We employ eight metrics to evaluate the quality of the fusion results, which are entropy (EN) [49], visual information fidelity (VIF) [55], the sum of the correlations of differences (SCD) [56], mean squared error (MSE), average gradient (AG)[57], correlation coefficient (CC), edge retentiveness ($Q^{AB/F}$) [58] and standard deviation (SD) [56]. EN measures the amount of information contained in an image based on the information theory. SD reflects the contrast of the fusion image. VIF measures the information fidelity of the fused image. MSE is the mean error value between the fusion image between source images. AG quantifies the gradient information. $Q^{AB/F}$ measures the amount of edge information transferred from source images to the fused image. SCD and CC compute the degree of correlation degree between source images and the fusion result. Furthermore, as described in Section 3.3, for our small-size image pairs, deep features of the VGG-16 network contain poor intensity and detail information, so we investigate the performance of different depth VGG-16 features used in the loss function. Lastly, we visualize the spatial attention maps in the dual attention fusion block to validate the ability of information integration.
@@
Figure 5: Qualitative comparison of our method with 7 state-of-the-art models on six infrared and visible image pairs in the RoadScene dataset. The first and second rows are visible and infrared images, respectively. From the third to the last rows are fusion results of DenseFuse[12], DRF[38], RFN-Nest[23], GANMcC[15], STDFusionNet[40], UMF-CMGR[31], YDTR[30] and our MDA.
IV-A Results on TNO Dataset
@@
IV-A1 Qualitative Evaluation
Fig. 4 exhibits qualitative comparison results of different models on the TNO dataset. The top two rows and the last row represent the visible images, infrared images and fusion results of our method, respectively. From the third to the seventh row are fusion results of DenseFuse, DRF, RFN-Nest, GANMcC, STDFusionNet, UMF-CMGR and YDTR. First of all, as the purpose of MDA is to effectively measure and integrate the complementary information, we mainly follow whether both the thermal radiation information of the infrared image and the detailed information of the visible image are effectively retained. Secondly, style loss can keep the fusion result maintain the same pixel intensity distribution as the infrared image. Lastly, the application of the multi-scale framework and patch-wise loss function drives the model to measure complementary information not only for the whole image but also for each local region.
Visual inspection shows that our results can preserve the high contrast of pixel intensity like the brighter targets in infrared images or illuminated regions in visible images. Meanwhile, our results preserve rich texture details from visible images and regions with dramatic pixel intensity changes in infrared images, such as the first column of Fig. 4. As is shown in the red box of this image pair, a person standing near the door with a higher temperature is highlighted well as the infrared image, and tree branches in the visible image are also retained. In addition, our results have fewer artifacts at the edge of regions where infrared and visible pixel intensity distributions collide compared with other methods, as can be seen from the reb boxes of the fourth, fifth and sixth columns in Fig. 4, where the edge region of people and background keeps sharp.
The fusion result of the sixth and last column in Fig. 4 illustrates the superiority of our multiscale framework and patch-wise loss function, where both thermal and detail information in the region of running person is well preserved, such as the junction of clothes and pants. Other methods like GANMcC prefer to retain the global pixel intensity of the infrared image, while the fusion result of RFN-Nest and STDFusionNet in the local region lacks thermal saliency. Thus, our method is able to measure and integrate vital information both in global and local areas.
TABLE I: Quantitative comparison of our method with 7 state-of-the-art models on the TNO and RoadScene dataset. The best performance are shown in bold, and the second and third best performance are shown in red and blue, respectively.
Method TNO Dataset RoadScene Dataset
@@ IV-A1 Qualitative Evaluation
DenseFuse 6.402 0.680 1.528 0.031 2.252 0.545 0.259 8.996 6.839 0.668 1.360 0.027 3.742 0.659 0.363 9.573
DRF 6.761 0.674 1.299 0.070 2.406 0.445 0.220 9.998 7.324 0.554 0.998 0.066 3.289 0.482 0.243 10.495
RFN-Nest 7.075 0.799 1.795 0.045 2.924 0.544 0.277 9.790 7.300 0.660 1.681 0.043 3.228 0.645 0.281 10.109
GANMcC 6.808 0.712 1.676 0.054 2.472 0.549 0.235 9.372 7.011 0.637 1.445 0.072 3.767 0.646 0.313 9.916
STDFusionNet 6.394 0.832 1.196 0.102 3.412 0.419 0.247 8.546 6.584 0.724 0.801 0.150 3.876 0.528 0.296 8.594
UMF-CMGR 6.651 0.749 1.636 0.032 3.154 0.539 0.402 9.222 7.047 0.733 1.461 0.029 4.125 0.644 0.484 9.821
YDTR 6.527 0.747 1.515 0.044 2.543 0.513 0.317 9.274 6.915 0.714 1.416 0.039 4.084 0.643 0.431 10.256
Ours 7.129 0.861 1.858 0.044 3.420 0.550 0.356 9.761 7.370 0.757 1.790 0.037 4.472 0.662 0.401 10.356
@@
TABLE II: Quantitative results of ablation study on the TNO dataset. Bold indicates the best performance.
@@
summation 7.132 0.848 1.621 0.059 2.723 0.469 0.242 9.270
concatenation 7.111 0.829 1.814 0.047 3.148 0.541 0.315 10.075
$w/o$ adaptive $L_{image}$ 7.094 0.771 1.781 0.044 2.211 0.512 0.167 9.691
$w/o$ adaptive $L_{patch}$ 7.164 0.722 1.769 0.044 1.949 0.523 0.132 9.746
$L_{pixel}$: pooling 1-3 6.933 0.786 1.556 0.052 3.429 0.468 0.175 9.185
$L_{pixel}$: pooling 1-4 7.134 0.756 1.768 0.043 2.089 0.525 0.145 9.703
Ours 7.129 0.861 1.858 0.044 3.420 0.550 0.356 9.761
IV-A2 Quantitative Evaluation
Quantitative comparison results listed in Table I show that our method achieves comparable results on the TNO dataset. From Table I we can see that our method MDA achieves the largest average value in VIF, SCD, CC and $Q^{AB/F}$, and obtain comparable statistics value on EN, MSE and AG. For EN and SD, images with higher contrast usually achieve better performance in these two metrics, such as DRF. However, our MDA expects to preserve pixel intensity and texture detail simultaneously, and its results retain comparable contrast while achieving more realistic edges, which satisfy human visual perception better. SCD, CC and $Q^{AB/F}$ show that our results maintain rich edge information transferred from the input image pair. These eight statistics values demonstrate that our results preserve the largest amount of complementary information transferred from the input image pairs with fewer artifacts, which satisfies the human visual system.
IV-B Results on RoadScene Dataset
IV-B1 Qualitative Evaluation
Fig. 5 exhibits qualitative comparison results of different models on the RoadScene dataset. We select 7 image pairs to illustrate fusion performance comparison for the self-driving scene, including day and night. Our fusion results exhibit minimal artifacts compared with the other seven state-of-the-art methods, such as the car with a clear structure in the red box of the first and third columns. None of the pedestrian targets are salient enough in visible images of the last four columns, and almost all the fusion results succeed in focusing on the pedestrian targets. Among them, the fusion performance of UMF-CMGR, YDTR and our MDA keep the most similar style to the infrared image. However, as with results presented in the TNO dataset, our results retain both thermal and detail information in the red box region, which indicates that our method is capable of adaptively measuring and integrating complementary information for global and local areas.
IV-B2 Quantitative Evaluation
Quantitative comparison results are listed in Table I, and our results are the best for VIF, SCD, CC and $Q^{AB/F}$, as the quantitative results in the TNO dataset, which indicates that our fusion results adaptively preserve the edge and detail information of input image pairs. Specifically, the mean value of $Q^{AB/F}$ in TNO and RoadScene datasets is much higher than in the other models, corresponding to the thermal targets’ sharp edge in the fusion images of the quantitative experiment. For EN, AG and SD, our results are not optimal but not too far from the best results, which can be considered as a trade-off of adding complementary information from the visible image to the higher contrast infrared image. In summary, our fusion results generate rich texture detail information meanwhile the preservation of saliency thermal targets is comparable.
IV-C Ablation Study
@@
In our study, we measure the complementary information via the statistical methods and features extracted by pre-trained VGG-16. In this section, we perform the ablation study to validate the effectiveness of the complementary information measurement and information aggregation structure in the network.
IV-C1 Evaluation of the fusion block
@@
As we discussed in the previous section, dual attention fusion block combines spatial and channel attention modules to integrate complementary information and suppress redundant information. Specifically, the spatial attention module exploits the inter-spatial relationship between the infrared and visible image. Based on the above consideration, we visualize the spatial attention map ${SM}_{ir}$, ${SM}_{vis}$ before multiplying with $F_{ir}$, $F_{vis}$ to validate the spatial attention map is capable of focusing on regions contains vital information. In Fig. 6, we can see that spatial attention maps ${SM}_{ir}$ and its corresponding inputs $F_{ir}$ are highly correlated. In this example image, the vital information for infrared feature maps $F_{ir}$ are regions including person and door, which are not significant in the visible feature maps $F_{vis}$. However, these regions in visible attention maps are highlighted, which means the information between two modalities has been effectively integrated via the dual attention fusion block.
@@
Figure 7: Illustration of fusion strategy results. The first row indicates the input infrared and visible image pair, and three fusion results in the second row are our MDA, replacing the dual attention fusion block with the elementary summation and concatenation, respectively.
@@
Figure 9: Results of different layers of VGG-16 feature maps used to estimate adaptive weight coefficients. The first row indicates the input infrared and visible image pair, and the second row is our MDA model ( only use feature maps before the first and the second pooling layers of VGG-16 to estimate adaptive weight coefficient of loss function ), use feature maps before the 1,2,3 pooling layers of VGG-16 and use feature maps before the 1,2,3,4 pooling layers of VGG-16, respectively.
IV-C2 Evaluation of the adaptive weight
To validate the effectiveness of our adaptive weight in measuring the complementary information, we perform two ablation experiments for fixed weight and different VGG-16 feature depths in weight generation. We perform the first experiment where $\omega^{ir}$ and $\omega^{vis}$ of $L_{image}$ or $L_{patch}$ are set to 0.5, respectively. The qualitative comparison results on the TNO dataset are shown in Fig. 8. The first fusion result is generated by our MDA method, which adopts adaptive weight coefficient in both $L_{image}$ and $L_{patch}$, and the last two fusion results are only fixing the weight of $L_{image}$ and $L_{patch}$, respectively. We can see that without the adaptive weight generated by the method proposed in this paper, fusion results of both fixed weight lack accurate detailed information.

### v2

@@ 4.1 Implementation
In this section, we verify the effectiveness of our method through qualitative and quantitative evaluations by comparing it with several state-of-the-art methods on the TNO, RoadScene and $M^{3}FD$ datasets. These comparative methods include the CNN-based method, GAN-based methods, AE-based and Transformer-based methods, which are DenseFuse [16], RFN-Nest [28], GANMcC [19], UMF-CMGR[39], SwinFusion[36], LRRNet[37] and FAFusion[38]. Qualitiative evaluation is based on the human visual system, hoping to highlight the thermal target throughout the scene while preserving texture detail. Quantitative evaluation access fusion image quality the way of several statistics and visual metrics. We employ eight metrics to evaluate the quality of the fusion results, which are entropy (EN) [58], visual information fidelity (VIF) [65], the sum of the correlations of differences (SCD) [66], mean squared error (MSE), average gradient (AG)[67], correlation coefficient (CC), edge retentiveness ($Q^{AB/F}$) [68] and standard deviation (SD) [66]. EN measures the amount of information contained in an image based on the information theory. SD reflects the contrast of the fusion image. VIF measures the information fidelity of the fused image. MSE is the mean error value between the fusion image between source images. AG quantifies the gradient information. $Q^{AB/F}$ measures the amount of edge information transferred from source images to the fused image. SCD and CC compute the degree of correlation degree between source images and the fusion result. Furthermore, as described in Section 3.3, for our small-size image pairs, deep features of the VGG-16 network contain poor intensity and detail information, so we investigate the performance of different depth VGG-16 features used in the loss function. Lastly, we visualize the spatial attention maps in the dual attention fusion block to validate the ability of information integration.
@@
Figure 4: Qualitative comparison of our method with 7 state-of-the-art models on seven infrared and visible image pairs in the TNO dataset. The first and second rows are visible and infrared images, respectively. From the third to the last rows are fusion results of DenseFuse[16], RFN-Nest[28], GANMcC[19], UMF-CMGR[39], SwinFusion[36], LRRNet[37], FAFusion[38] and our MDA.
Figure 5: Qualitative comparison of our method with 7 state-of-the-art models on seven infrared and visible image pairs in the RoadScene dataset. The first and second rows are visible and infrared images, respectively. From the third to the last rows are fusion results of DenseFuse[16], RFN-Nest[28], GANMcC[19], UMF-CMGR[39], SwinFusion[36], LRRNet[37], FAFusion[38] and our MDA.
4.2 Results on TNO Dataset
@@
4.2.1 Qualitative Evaluation
Fig. 4 exhibits qualitative comparison results of different models on the TNO dataset. The top two rows and the last row represent the visible images, infrared images and fusion results of our method, respectively. From the third to the ninth row are fusion results of DenseFuse, RFN-Nest, GANMcC, UMF-CMGR, SwinFusion, LRRNet and FAFusion. First of all, as the purpose of MDA is to effectively measure and integrate the complementary information, we mainly follow whether both the thermal radiation information of the infrared image and the detailed information of the visible image are effectively retained. Secondly, style loss can keep the fusion result maintaining the same pixel intensity distribution as the infrared image. Lastly, the application of the multi-scale framework and patch-wise loss function drives the model to measure complementary information not only for the whole image but also for each local region.
Visual inspection shows that our results can preserve the high contrast of pixel intensity like the brighter targets in infrared images or illuminated regions in visible images. Meanwhile, our results preserve rich texture details from visible images and regions with dramatic pixel intensity changes in infrared images, such as the first column of Fig. 4. As is shown in the red box of this image pair, a person standing near the door with a higher temperature is highlighted well as the infrared image, and the chair which is only can be seen in the visible image is also retained. Our goal is to adaptively measure and integrate complementary information from both modalities, therefore we can see from the fourth column that the branch with complex texture of the infrared image is preserved well, and so does the crawling person with higher pixel intensity. In addition, our results have fewer artifacts at the edge of regions where infrared and visible pixel intensity distributions collide compared with other methods, as can be seen from the reb boxes of the fourth, fifth and sixth columns in Fig. 4, where the edge region of people and background keeps sharp.
The fusion result of the sixth and last column in Fig. 4 illustrates the superiority of our multiscale framework and patch-wise loss function, where both thermal and detail information in the region of running person is well preserved, such as the junction of clothes and pants. Other methods like GANMcC and SwinFusion prefer to retain the global pixel intensity of the infrared image, while the fusion result of RFN-Nest and LRRNet in the local region lacks thermal saliency. Thus, our method is able to measure and integrate vital information both in global and local areas.
Table 1: Quantitative comparison of our method with 7 state-of-the-art models on the TNO, RoadScene and $M^{3}FD$ dataset. The best performance are shown in bold, and the second and third best performance are shown in red and blue, respectively.
Method $TNO$ Dataset
@@ 4.2.1 Qualitative Evaluation
DenseFuse 6.984 0.826 1.829 0.043 3.475 0.545 0.271 9.679
RFN-Nest 7.077 0.825 1.829 0.047 2.682 0.549 0.271 9.726
GANMcC 6.967 0.760 1.727 0.047 2.694 0.546 0.271 9.726
UMF-CMGR 6.641 0.744 1.641 0.033 3.153 0.540 0.401 9.139
SwinFusion 6.924 0.911 1.675 0.059 4.207 0.501 0.493 9.577
LRRNet 7.127 0.835 1.604 0.051 4.029 0.476 0.347 9.457
FAFusion 6.819 0.634 1.179 0.060 5.041 0.364 0.346 9.362
Ours 7.111 0.850 1.850 0.044 3.524 0.550 0.357 9.735
Method $RoadScene$ Dataset
[…]
DenseFuse 6.839 0.668 1.360 0.027 3.742 0.659 0.363 9.573
RFN-Nest 7.300 0.660 1.681 0.043 3.228 0.645 0.281 10.109
GANMcC 7.011 0.637 1.445 0.072 3.767 0.646 0.313 9.916
UMF-CMGR 7.047 0.733 1.461 0.029 4.125 0.644 0.484 9.821
SwinFusion 6.922 0.780 1.591 0.060 4.491 0.624 0.455 10.474
LRRNet 7.135 0.701 1.569 0.068 4.703 0.620 0.349 10.070
FAFusion 6.991 0.559 0.798 0.061 5.694 0.464 0.371 10.088
Ours 7.370 0.757 1.790 0.037 4.472 0.662 0.401 10.356
Method $M^{3}FD$ Dataset
[…]
DenseFuse 6.698 0.762 1.695 0.032 3.302 0.585 0.503 9.021
RFN-Nest 6.863 0.773 1.726 0.034 2.855 0.572 0.405 9.232
GANMcC 6.624 0.664 1.576 0.037 2.412 0.571 0.267 9.002
UMF-CMGR 6.699 0.709 1.570 0.032 2.929 0.546 0.397 9.144
SwinFusion 6.790 0.877 1.562 0.060 4.615 0.521 0.597 9.859
LRRNet 6.437 0.719 1.463 0.039 3.600 0.542 0.497 9.307
FAFusion 6.895 0.689 1.254 0.051 5.799 0.439 0.422 9.692
Ours 6.909 0.816 1.778 0.033 3.602 0.585 0.487 9.737
4.2.2 Quantitative Evaluation
Quantitative comparison results listed in Table 1 show that our method achieves comparable results on the TNO dataset. From Table 1 we can see that our method MDA achieves the largest average value in SCD, CC and SD, and obtain comparable statistics values on EN, VIF, MSE, AG and $Q^{AB/F}$. For EN and SD, images with higher contrast usually achieve better performance, such as RFN-Nest and GANMcC hihglight the human in the fourth and fifth columns. However, our MDA expects to preserve pixel intensity and texture detail simultaneously, and its results retain comparable contrast while achieving more realistic edges, which satisfy human visual perception better. SCD, CC and AG show that our results maintain rich edge information transferred from the input image pair. These eight statistics values demonstrate that our results preserve the largest amount of complementary information transferred from the input image pairs with fewer artifacts, which satisfies the human visual system.
Figure 6: Qualitative comparison of our method with 7 state-of-the-art models on seven infrared and visible image pairs in the $M^{3}FD$ dataset. The first and second rows are visible and infrared images, respectively. From the third to the last rows are fusion results of DenseFuse[16], RFN-Nest[28], GANMcC[19], UMF-CMGR[39], SwinFusion[36], LRRNet[37], FAFusion[38] and our MDA.
4.3 Results on RoadScene Dataset
4.3.1 Qualitative Evaluation
Fig. 5 exhibits qualitative comparison results of different models on the RoadScene dataset. We select 7 image pairs to illustrate fusion performance for the self-driving scene, including day and night. Our fusion results exhibit minimal artifacts compared with the other seven state-of-the-art methods, such as the car with a clear structure in the red box of the first and third columns. None of the pedestrian targets are salient enough in visible images of the last four columns, and almost all the fusion results succeed in focusing on the pedestrian targets. Among them, the fusion performance of UMF-CMGR and our MDA keep the most similar style to the infrared image. However, as with results presented in the TNO dataset, our results retain both thermal and detail information in the red box region, which indicates that our method is capable of adaptively measuring and integrating complementary information for global and local areas. Additionally, the texture of the sky and branches in the third and fourth columns are preserved, meanwhile the thermal target maintains a high contrast with the surrounding environment in our MDA method, suggesting that the intensity and detail information from the infrared images are adaptively measured.
4.3.2 Quantitative Evaluation
Quantitative comparison results are listed in Table 1, and our results are the best for EN, SCD and CC, as the quantitative results in the TNO dataset, which indicates that our fusion results adaptively preserve the edge and detail information of input image pairs. Specifically, the mean value of SCD and CC in TNO and RoadScene datasets is higher than other models, corresponding to the richest transferred information from the input image pairs. For VIF and SD, our results are not optimal but not too far from the best results, which can be considered as a trade-off of adding complementary information from the visible image to the higher contrast infrared image. In summary, our fusion results generate rich texture detail information meanwhile the preservation of saliency thermal targets is comparable.
4.4 Results on $M^{3}FD$ Dataset
We have also tested our method on the $M^{3}FD$ dataset, which consists of high-resolution color visible and grayscale infrared image pairs, to explore the fusion performance in night and bad weather scenarios.
4.4.1 Qualitative Evaluation
Fig. 6 exhibits qualitative comparison results of different models on the $M^{3}FD$ dataset. We select 7 image pairs to exhibit fusion performance comparison for poor visual conditions, like rainy nights and smog. Due to bad weather conditions and smog, the thermal targets and texture information of visible images are obscured by the smog. Almost all the methods are able to highlight the thermal target in the infrared images, but some of them fail to preserve detail information behind the smog, such as the SwinFusion. Our method preserves the high contrast between the target and background of the infrared image, while the complex texture only can be seen in the infrared image is also preserved, such as the mountain and buildings behind the smog in the first, fifth and last columns.
The fusion results of our MDA keep a similar intensity style with infrared images, such as the fused images from the second to fourth columns. Visible images are full of fog, which causes halo effect around the light and low contrast. Our fusion results preserve high contrast from the infrared images and halo effect is reduced, like the sharp traffic light in the third column.
4.4.2 Quantitative Evaluation
Quantitative comparison results are listed in Table 1, and our results are the best for EN, SCD and CC, meanwhile the VIF, MSE, AG and SD also achieve comparative results. The largest value of CC and SCD indicate that our fusion results are highly correlated with the original image pairs, and the second largest SD and AG values of our method indicate that our method keeps comparative intensity contrast and texture detail information simultaneously.
@@ 4.4.2 Quantitative Evaluation
Table 2: Quantitative results of ablation study on the TNO dataset. The best performance are shown in bold, and the second and third best performance are shown in red and blue, respectively.
@@
summation 7.125 0.845 1.617 0.060 2.728 0.467 0.242 9.271
concatenation 7.082 0.820 1.816 0.047 3.159 0.543 0.316 10.072
$w/o$ adaptive $L_{image}$ 7.080 0.763 1.774 0.043 2.231 0.529 0.167 9.676
$w/o$ adaptive $L_{patch}$ 7.149 0.712 1.760 0.044 1.962 0.521 0.132 9.737
$L_{pixel}$: pooling 1-3 6.940 0.783 1.559 0.054 3.476 0.467 0.176 9.194
$L_{pixel}$: pooling 1-4 7.120 0.746 1.760 0.043 2.105 0.522 0.146 9.691
Ours 7.111 0.850 1.850 0.044 3.524 0.550 0.357 9.735
4.5 Ablation Study
@@
In our study, we measure the complementary information via the statistical methods and features extracted by pre-trained VGG-16. In this section, we perform the ablation study to validate the effectiveness of the complementary information measurement, information aggregation structure and hyperparameters in the network.
4.5.1 Evaluation of the fusion block
@@
As we discussed in the previous section, dual attention fusion block combines spatial and channel attention modules to integrate complementary information and suppress redundant information. Specifically, the spatial attention module exploits the inter-spatial relationship between the infrared and visible image. Based on the above consideration, we visualize the spatial attention map ${SM}_{ir}$, ${SM}_{vis}$ before multiplying with $F_{ir}$, $F_{vis}$ to validate the spatial attention map is capable of focusing on regions contains vital information. In Fig. 7, we can see that spatial attention maps ${SM}_{ir}$ and its corresponding inputs $F_{ir}$ are highly correlated. In this example image, the vital information for infrared feature maps $F_{ir}$ are regions including person and door, which are not significant in the visible feature maps $F_{vis}$. However, these regions in visible attention maps are highlighted, which means the information between two modalities has been effectively interacted with and integrated via the dual attention fusion block. Likewise, the infrared attention maps contain rich detail texture information in the region of plants, where the input infrared image has poor texture in the same region.
@@
Figure 8: Illustration of fusion strategy results. The first row indicates the input infrared and visible image pair, and three fusion results in the second row are our MDA, replacing fusion block with the summation and concatenation, respectively.
@@
Figure 10: Results of different layers of VGG-16 feature maps used to estimate adaptive weight coefficients. The first row indicates the input infrared and visible image pair, and the second row is our MDA model, using feature maps before the 1,2,3 pooling layers of VGG-16 and feature maps before the 1,2,3,4 pooling layers of VGG-16, respectively.
4.5.2 Evaluation of the adaptive weight
To validate the effectiveness of our adaptive weight in measuring the complementary information, we perform two ablation experiments for fixed weight and different VGG-16 feature depths in weight generation. We perform the first experiment where $\omega^{ir}$ and $\omega^{vis}$ of $L_{image}$ or $L_{patch}$ are set to 0.5, respectively. The qualitative comparison results on the TNO dataset are shown in Fig. 9. The first fusion result is generated by our MDA method, which adopts adaptive weight coefficient in both $L_{image}$ and $L_{patch}$, and the last two fusion results are only fixing the weight of $L_{image}$ and $L_{patch}$, respectively. We can see that without the adaptive weight generated by the method proposed in this paper, fusion results of both fixed weights lack accurate detailed information.